OpenStudy (anonymous):

How can you figure out a tangent without using a calculator? Like the tangent of -pi.

OpenStudy (anonymous):

You really can't for some angles without a calculator. However, for pi, tan(-pi) = - tan(pi) = 0 (sin(pi) = 0). For some standard angles like pi or its multiples, these values can be calculated.

OpenStudy (anonymous):

Okay so you can only do some angles without a calculator. Can you do 3(1-cosx) without a calculator?

OpenStudy (anonymous):

Again, it depends on what x is. If x is some multiple of pi, the value can be calculated easily. So. 3(1 - cospi) = 3(1 -(-1)) = 6.